# Translators Asymptotic to Planes

### v1

@@ 2. Preliminaries
For each $\Sigma_{i}$, up to an ambient isometry we may assume $\Pi_{i}=\{x_{1}=0\}$ and $p_{i}=0$. We introduce polar coordinates for $\Pi_{i}$, conflating $(0,r\cos\theta,r\sin\theta)$ with $(r,\theta)$. We may then express points in $\Sigma_{i}$ as $(u_{i}(r,\theta),r,\theta)$. For all sufficiently large $R$ we define map
@@
$\lim_{R\to\infty}\int_{\Gamma_{R,i}}\langle\eta,e_{3}^{\top}\rangle=0$
@@
Our aim now is to prove that any end which is asymptotic to a vertical plane automatically satisfies (2). This will make use of the following interior gradient estimate, which is obtained by a standard application of Bernstein’s technique. [^1] Møller pointed out to us that the same estimate can be obtained by the different arguments contained in [2, Proposition 3.10] and a standard interpolation inequality.

### v2

@@ 1. Introduction
We note that since $\Sigma$ is not assumed to be embedded in Theorem 1.1, its asymptotic vertical planes need not be parallel.
@@
Remark 1.3.
After originally posting this article we learned of the work [6], by Neves and Tian, concerning the classification of Lagrangian translators. Some elements of our proof are similar to theirs—in particular, they derive the same flux formula for the total mean curvature and utilise it to prove flatness in their setting.
Acknowledgements. We are grateful to Felix Schulze for bringing to our attention the reference [6].
@@ 2. Preliminaries
For each $\Sigma_{i}$, up to an ambient isometry we may assume $\Pi_{i}=\{x_{1}=0\}$ and $p_{i}=0$. We introduce polar coordinates for $\Pi_{i}$, conflating $(0,r\cos\theta,r\sin\theta)$ with $(r,\theta)$. We may then express points in $\Sigma_{i}$ as $(u_{i}(r,\theta),r,\theta)$. For all sufficiently large $R$ we define a map
@@
$\lim_{R\to\infty}\int_{\Gamma_{i,R}}\langle\eta,e_{3}^{\top}\rangle=0$
@@
Our aim now is to prove that any end which is asymptotic to a vertical plane automatically satisfies (2). This will make use of the following interior gradient estimate, which is obtained by a standard application of Bernstein’s technique. [^1] We note that Gama, Martín and Møller have independently obtained the same estimate by different arguments—it is implied by [2, Proposition 3.10] via a standard interpolation inequality.